OpenStudy (anonymous):

Essay. Show all work. y^4+y+1/y^2+3 and 9a^2-100b^2/3a-10b

OpenStudy (anonymous):

What is the complete question? :(

OpenStudy (anonymous):

Essay: Show all work. y^4+y+1/y^2+3 is the first one. The second question is 9a^2-100b^2/3a-10b

OpenStudy (shayaan_mustafa):

Hi fthrogers :) Your question is still unclear to me..

OpenStudy (anonymous):

The first set of numbers are one problem and the second set is another problem.

OpenStudy (shayaan_mustafa):

Does your question looks like this?\[\large \frac{y^4+y+1}{y^2+3}\]and\[\large \frac{9a^2-100b^2}{3a-10b}\]

OpenStudy (anonymous):

yes

OpenStudy (shayaan_mustafa):

and what to do with it?

OpenStudy (anonymous):

Find the quotient

OpenStudy (anonymous):

Sorry, I didnt type the whole sentence.

OpenStudy (shayaan_mustafa):

oh I see. quotient for the first part is going to be long.

OpenStudy (shayaan_mustafa):

\[\large \frac{9a^2-100b^2}{3a-10b}\]use formula (a^2-b^2)=(a+b)(a-b) for numerator.\[\large \frac{\cancel{(3a-10b)}(3a+10b)}{\cancel{3a-10b}}\]\[\large 3a+10b\]It is your quotient for the second part.

OpenStudy (anonymous):

Thank you

OpenStudy (shayaan_mustafa):

can you do it for the first part?

OpenStudy (anonymous):

No, I can try, but Im really unsure.

OpenStudy (anonymous):

Is there anyway that you can help me?

OpenStudy (shayaan_mustafa):

yes i can help you friend. Do you know how to use synthetic division?

OpenStudy (anonymous):

No, I am really not good at algebra but I am trying to figure out how to start figuring the problem, so maybe I can understand it.

OpenStudy (shayaan_mustafa):

Wait please . I will solve it for you.

OpenStudy (anonymous):

Thank you so much, I really need to learn to understand so that I can do these kinds of problems. I have college algebra next semester and I am terrified.

OpenStudy (shayaan_mustafa):

Sorry friend. my mouse is flickering regularly therefore i am unable to solve the first one. i apologize for it. and first one as you can see that it is just substitute formula and cancel. while second one we have to do long division.

OpenStudy (anonymous):

That is okay, i will work on it to see what I can come up with..